OpenStudy (anonymous):

Simplify (4xy)(2x^2y)(3xy)^3.

OpenStudy (anonymous):

Is it?? \[\large (4xy)(2x^y)(3xy)^3\]

OpenStudy (anonymous):

So, you cleverly changed the question.. :) That's good..

OpenStudy (anonymous):

Solve one by one.. Use this concept: \[\large (mn)^a = m^a \cdot n^a\] Exponent outside brackets becomes exponents of individual variables inside the brackets. \[\large a^m \cdot a^n = a^{m+n}\] When base is equal, then exponents get added..

OpenStudy (anonymous):

like I tried working it out but I got confused

OpenStudy (anonymous):

Tell me where you got confused??

OpenStudy (anonymous):

I'm confused with the exponents

OpenStudy (anonymous):

Sorry, I read the question incorrectly.. You can eliminate two options by just multiplying the constants in the question.. \[4 \times 2 \times 3^3 = ??\]

OpenStudy (anonymous):

this is confusing :/